Wednesday, March 15, 2000 (DOY 068/19:00:00 to DOY 075/19:00:00 UTC)

Launch / Days since Launch = Nov 7, 1996 / 1227 days

Start of Mapping / Days since Start of Mapping = April 1, 1999 / 349 days
Total Mapping Orbits =

Total Orbits = 6165

Recent Events:

The spacecraft continues to operate nominally in performing the beta supplement daily recording and
transmission of science data. The mm020 sequence executed successfully from 00-63 (3/3/00)
through 00-72 (3/12/00). The mm021 sequence also successfully executed from 00-63 (3/13/00)
through 00-74 (3/14/00). The mm022 sequence was successfully uplinked on 00-73 (3/13/00) and
began execution nominally on 00-74 (3/14/00).

Upon completion of the recent four day fixed-HGA mapping campaign performed last week, which
resulted in higher than expected MOLA temperatures, the spacecraft team began looking at ways to
fix the temperature problem. Additionally in assessing the first fixed-HGA campaign, it was noted that
the actual transitions into and out of fixed-HGA mode of operations from the beta supplement mode
of operations, were more complicated than anticipated and resulted in a larger than expected loss of
science data from the other instruments. The Spacecraft Team presented at the 3/14 Project Mission
Planning meeting alternatives to and options for fixing the current fixed-HGA operations. Fixed-HGA
operations were to be performed 4-8 days per month for the express purpose of acquiring Radio
Science Earth occultation egress measurements, which are prohibited during normal beta supplement
mapping operations due to HGA gimbal constraints. From the Mission Planning meeting the favored
option was to eliminate the fixed-HGA operations and perform special slews for a specified number of
orbits to provide the desired longitudinal planetary coverage for the radio science occultation egress
data, similar to the implementation of the MOLA off-nadir polar observations. Final details are in work
and will be presented to the Project Change Board next Tuesday (3/21)

Spacecraft Health:

All subsystems are reporting nominal health. Battery 2 was successfully commanded back to VT-2
after transitioning back to beta supplement operations. The red alarm high limit for the MOLA
temperature has been decreased by 3 degrees (from 35°C to 32°C) after a meeting with the MOLA


There have been 20 uplinks to the spacecraft during the last week, including new star catalogs and
ephemeris files, instrument command loads, and the mm021 and mm022 sequences. Two uplinks were
missed after the final transition from the fixed-HGA Earth pointed attitude back to nadir pointed beta
supplement operations due to a large change in the transponder temperature. The uplink sweeps for
the beta supplement orbits are tailored specially for the very short uplink windows and are very
sensitive to changes in the MOT temperature. The missing uplinks were successfully re-radiated. Total
command files radiated to the spacecraft since launch is 4515.

Upcoming Events:

Planning has begun for a focus calibration of the MOC to be performed in late March.